OpenStudy (anonymous):

Which of the following options is an equivalent function to f(x) = 2(5)2x? f(x) = 50x f(x) = 100x f(x) = 2(25)x f(x) = 4(25)x

OpenStudy (jdoe0001):

hint: \(\bf (a^n)^m\implies (a^m)^n\implies (a^{n\cdot m})\qquad thus \\ \quad \\ 2(5)^{2x}\implies 2(5^2)^x\implies 2(5^x)^2\implies 2(5^{2x})\)

OpenStudy (jdoe0001):

think about it \(\bf 2(5)^{2x}\implies \underline{2(5^2)^x}\implies 2(5^x)^2\implies 2(5^{2x})\)

OpenStudy (anonymous):

f(x) = 100x?

OpenStudy (anonymous):

@jdoe0001

OpenStudy (jdoe0001):

haemmm nope the 2 is not affected by the exponent, since the 2 is not inside the parentheses

OpenStudy (jdoe0001):

notice \(\bf \bf 2(5)^{2x}\implies {\color{brown}{ 2(5^2)^x }}\implies 2(5^x)^2\implies 2(5^{2x}) \)

OpenStudy (anonymous):

it will be \[2(5)^{2x}\] this will lead to \[2([5^2]^x)\] so this leads to f(x)=\[2(25)^{x}\]

OpenStudy (anonymous):

f(x)=2(25)x

OpenStudy (fairygirl0404):

None 2(5)2x = 20x If you meant 2(5)^2x, then 2(5)^2x = 2(25)x = 50x.